OpenStudy (anonymous):

1.6≥x over -4 (Show Work)

OpenStudy (anonymous):

x≥-6.4 and the solutuon is x≥-32/5

OpenStudy (anonymous):

1.6≥x 1.6*(10/10)>=x 16/10>=x 8/5>=x x<=8/5